OpenStudy (openstudyrocks5*):

Explain why the Ratios 25/45 & 15/27 are equivalent

OpenStudy (pawanyadav):

divide 25/45 by3 Both in numerator and denominator

OpenStudy (pawanyadav):

Sorry by 5

OpenStudy (pawanyadav):

And 15/27 By 3

OpenStudy (pawanyadav):

Then you will get your answer

OpenStudy (openstudyrocks5*):

@Pawanyadav So sorry I didn't respond or anything, I was in the bathroom feeling sick ;_;

OpenStudy (openstudyrocks5*):

@Pawanyadav Ok I did it and got 1/9 and 1/5, That's Equal? or did I do it wrong

OpenStudy (pawanyadav):

Do again 1st by 5

OpenStudy (pawanyadav):

25÷5/45÷5=?

OpenStudy (openstudyrocks5*):

Hmm I must be doing something wrong, Now I got 1 and 9 xD Damn... Could you show me EXACTLY how to divide it and stuff

OpenStudy (openstudyrocks5*):

@Pawanyadav

OpenStudy (openstudyrocks5*):

Oh sorry my thing glitched. So...

OpenStudy (openstudyrocks5*):

So.. May I ask what 5/9 does with it being Equivilent.. Just asking not being harmfully sarcastic.

OpenStudy (pawanyadav):

What is 15÷3=? And 27÷3=?

OpenStudy (openstudyrocks5*):

@Pawanyadav Haha nvm I get it, Thanks Pawan, sorry about the confusion, im a piece of work xD

OpenStudy (pawanyadav):

Thanks you got it